OpenStudy (anonymous):

Convert the equation from rectangular to polar form 9xy = 1

OpenStudy (phi):

you can use x = r cosA y = r sin A

OpenStudy (jdoe0001):

\(\bf x=rcos(\theta)\qquad y=rsin(\theta) \\ \quad \\ \quad \\ 9xy = 1\implies 9[rcos(\theta)][rsin(\theta)]=1\implies [rcos(\theta)][rsin(\theta)]=\cfrac{1}{9}\) take common factor on the left-side and solve for "r"

OpenStudy (anonymous):

So would it be \[ r = 1 / 9\cos(\theta)\sin(\theta)\]

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Thank you!

OpenStudy (jdoe0001):

yw